OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y - y1 = 4(x - x1) for your point (x1, y1)

OpenStudy (anonymous):

Can you see which is correct?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

:)

hero (hero):

General Form of point-slope formula: y - y1 = m(x-x1) After inputting point (-3,-1), you'll end up with: y + 1 = 4(x + 3)